OpenStudy (anonymous):

How do I solve log(base 2)8^4?

OpenStudy (jdoe0001):

\(\large \bf {{\color{blue}{ log_ab=\square \implies a^\square =b}}\\ \quad \\ \quad \\ log_28^4=\square \implies 2^\square =8^4\qquad \qquad {\color{blue}{ 8\implies 2^3}}\quad thus\\ \quad \\ 2^\square =8^4\implies 2^\square =(2^3)^4}\)

OpenStudy (anonymous):

What are the boxes? like it says log(base a)b= a box?

OpenStudy (jdoe0001):

yes, the value you're looking for, the exponent for the base of "2"

OpenStudy (jdoe0001):

\(\bf (2^3)^4\implies 2^{\color{red}{ \square}} ?\)

OpenStudy (anonymous):

Oh lol okay. Thank you! It worked :)

OpenStudy (jdoe0001):

yw